# Human and LLM Biases in Hate Speech Annotations: A Socio-Demographic Analysis of Annotators and Targets

### v3

@@ Abstract
Warning: This paper contains examples that may be perceived as offensive or upsetting.
@@ Introduction
The proliferation of hateful and toxic speech on online social platforms is among the most severe threats to inclusive online spaces (Mathew et al. 2019). The exponential increase in user-generated content has made manual hate speech detection impractical, driving the development of automated systems (Garg et al. 2023). However, the effectiveness of automated detectors is heavily contingent upon the quality of their training data, which is typically human-labeled. Challenges arise as hate speech detection is inherently subjective: what one individual may perceive as offensive, another might consider acceptable or even normal. This subjectivity is shaped by the annotators’ backgrounds, experiences, and cultural contexts. As such, socio-demographic attributes such as age, gender, race, and others, can influence how annotators interpret and label content, possibly leading to biases (Davani et al. 2023). In turn, biases may seep into the automated detection systems trained on such data, causing severe consequences such as the unwarranted removal of non-hateful content or the disproportionate targeting of marginalized groups (Nogara et al. 2025).
@@
Large language models (LLMs) have recently extended the set of AI tools in support of platform administrators. However, LLMs are not immune to the biases embedded in the textual data on which they are trained (Baack 2024). Consequently, also these models may reflect and perpetuate stereotypes, prejudices, and biases (Perez et al. 2023). Despite the development of various techniques aimed at reducing LLM biases, this remains an outstanding challenge (Ouyang et al. 2022). Interestingly, the capacity of LLMs to produce subjective outputs can also serve beneficial purposes. LLMs can be personalized and deployed as actors in agent-based simulations of online human interactions (Törnberg et al. 2023), which represents a promising way to compensate for the limited availability of platform data in the post-API age. By accurately reproducing certain human biases, persona-based LLMs could lay the groundwork for the next generation of AI-driven social simulations. For these reasons, investigating biases in both human hate speech annotations and persona-based LLMs is a task of great theoretical and practical relevance (Tseng et al. 2024).
@@
In RQ1a, we hypothesize that identification with the target group (i.e., the annotator’s in-group) may heighten the annotator’s sensitivity to hate, leading them to flag such content more often (Sachdeva et al. 2022b). Conversely, when the target belongs to an out-group, the annotator’s perception of the hate may be less acute. RQ1b broadens the analysis by examining whether certain groups systematically label content differently than others, in relation to the targets of the hate. This analysis reveals possible disparities in how hate speech is perceived across diverse socio-demographic backgrounds.
@@
Bias in human hate speech annotations. The impact that biased datasets have on the reliability and fairness of automated hate speech detection systems sparked much research on the issue (Garg et al. 2023). Waseem and Hovy (2016) showed that cultural and socio-economic backgrounds lead to variability hate speech annotations. Inherent prejudices further skew labels, as human annotators may be more sensitive to hate speech against familiar groups while underestimating it against less familiar ones. Others found that socio-demographic factors such as age, gender, race and ethnicity, education, and English proficiency, influence annotation outcomes, with younger and minority annotators more likely to label content as hate speech (Al Kuwatly, Wich, and Groh 2020; Davani et al. 2023). Additionally, annotator beliefs and demographics can introduce further inconsistencies in crowdsourced annotations (Hettiachchi et al. 2023). Although much research has explored socio-demographic biases in hate speech annotations (Garg et al. 2023), most studies focused on a limited set of attributes, primarily race and ethnicity, gender, and age. Moreover, none investigated the interplay between annotator and target attributes.
@@
Demographic alignment of LLMs. LLM prompting has been used to experiment with persona-based and role-playing models, to test the alignment of LLM predictions with human opinions (Tseng et al. 2024). Studies involved replicating famous cognitive and social experiments (Aher, Arriaga, and Kalai 2023; Srivastava et al. 2023), as well as assessing agreement and personalization capabilities. Among the existing works, Beck et al. (2024) and Hu and Collier (2024) measured the impact of socio-demographic prompting on model performance, showing that it can enhance zero-shot learning in subjective tasks. Additional studies explored the personalities (e.g., Big5, MBTI) exhibited by LLMs when prompted to emulate certain human socio-psychological traits (Rao, Leung, and Miao 2023; La Cava, Costa, and Tagarelli 2024). Results showed marked differences between models and a moderate predisposition to personalization and psychological prompting. However, Santurkar et al. (2023) measured the alignment between human and LLM opinions across a wide array of demographic groups and topics, finding substantial misalignment. The misalignment persisted even after explicitly steering the LLMs towards particular groups. Similarly, Lee, An, and Thorne (2023) investigated the alignment of instruction-tuned LLMs with the disagreement among human annotators, finding that models fail to capture human disagreements. As briefly summarized, there are many contrasting results on the alignment of persona-based LLMs with human opinions. This gap in understanding underscores the need for further research to clarify the extent and conditions under which LLMs can mirror human perspectives.
@@
The dataset follows the FAIR principles. It is available on a prominent cloud storage service, [^1] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech (accessed: 15/09/2024) making it findable and accessible. It is interoperable in that it is released in a machine-readable format, and reusable due to the information contained in the original paper (Sachdeva et al. 2022a).
@@
Operationalization. Our definition of bias entails the systematic comparison between the labels assigned by human or LLM annotators with certain attributes, with the labels assigned by annotators without such attributes. We operationalize the comparison by extending the method of Wich et al. (2021) as follows. Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of annotators and $G=\{g_{1},\ldots,g_{m}\}$ be the set of targets. We first fix one socio-demographic attribute $\mathbf{t}$ and a specific value of that attribute $\mathbf{t}=v$ (e.g., ideology $=$ liberal). Next, we select all posts labeled by all annotators $A(\mathbf{t}=v)$ with the fixed attribute and value. Depending on the RQ to answer, we further filter the posts by only selecting those whose target $G(\mathbf{t^{\prime}}=v^{\prime})$ feature another, possibly different, attribute and value $\mathbf{t^{\prime}}=v^{\prime}$. Finally, we compare the labels assigned by the annotators $A(\mathbf{t}=v)$ to those of all other annotators $A(\mathbf{t}\neq v)$ on the same set of selected posts. The comparison is based on the confusion matrix $D$ in Figure 2. Rows in $D$ correspond to the labels assigned by the annotators $A(\mathbf{t}=v)$, while columns correspond to the labels assigned by $A(\mathbf{t}\neq v)$. The diagonal represents those instances that are labeled equally by the two groups. The lower triangle represents instances where the annotators $A(\mathbf{t}=v)$ overestimate the hate with respect to $A(\mathbf{t}\neq v)$. Conversely, the upper triangle represents instances where $A(\mathbf{t}=v)$ underestimate hate. We repeat the analysis $\forall\;\mathbf{t},v$. Algorithm A1 in the Appendix provides the pseudo-code implementation of the process used to compute the confusion matrix $D$.
Statistical significance. Our comprehensive approach to investigating bias across a large set of socio-demographic attributes results in a substantial number of comparisons. However, as reported in Appendix Table T1, certain minority groups of annotators and targets are underrepresented in the data. Consequently, some comparisons are based on limited data, and the observed labeling differences may lack statistical significance. We address this issue by conducting Mann-Whitney tests to assess the statistical significance of the differences between the distributions of labels assigned by any two groups of annotators being compared. In the remainder, we only report results about those comparisons where the labeling differences are statistically significant at the 0.05 level. Due to data imbalance, we were unable to obtain any significant result for a small set of particularly underrepresented groups, highlighted in gray in Table T1.
@@
Model selection. We prioritize open-source models over commercial options (e.g., GPT-4) when selecting the LLMs for our experiments. First, open-source models are freely available, which enhances transparency and reproducibility of our research. Additionally, many open-source models have comparable accuracy to that of leading commercial LLMs. [^2] https://lmarena.ai/?leaderboard (accessed: 15/09/2024) Open-source LLMs also require less computational resources, broadening accessibility and easing implementation. These choices make our results generalizable and broadly applicable. Driven by these considerations, we based our model selection on recent results on the performance of open-source LLMs, favoring high-performance and widely-used models (La Cava, Costa, and Tagarelli 2024). Our selection includes Llama-3-8B-Instruct (8B), [^3] https://shorturl.at/0xT0b (accessed: 15/09/2024) Phi-3-mini-4k-instruct (3.8B), [^4] https://shorturl.at/tOs9i (accessed: 15/09/2024) SOLAR-10.7B-Instruct-v1.0 (10B), [^5] https://shorturl.at/udLQZ (accessed: 15/09/2024) and Starling-LM-7B-alpha (7B). [^6] https://shorturl.at/kqqLq (accessed:15/09/2024)
@@
Figure 4: Intensity ($I$) and prevalence ($P$) of human biases in hate speech annotations, for all possible combinations of annotator (y axis) and target (x axis) socio-demographic attributes. Cell color corresponds to intensity while cell area is proportional to prevalence. Statistical significance is assessed at the 0.05 level with a Mann-Whitney test. Missing cells denote non-significant results. Red-colored cells mark statistically significant hate overestimations, while blue-colored ones denote significant underestimations. The magenta-colored top-left square corresponds to the socio-demographic attributes that are available for both the annotators and targets (age, gender, race, religion, sexuality), as thoroughly reported in Appendix Table T1. The dashed diagonal corresponds to in-group annotations (i.e., same annotator and target attribute). Attributes outside of the top-left square are only available either for annotators (education, ideology, income) or targets (disability, origin). Annotator and target attributes for which we did not obtain any significant result (e.g., ideology: extremely liberal) are not shown. When observed row-wise, the heatmap allows identifying groups of annotators with the tendency to systematically over- or underestimate hate. When observed column-wise, it allows identifying hate targets that are overall flagged more or less frequently. Single cells show bias intensity and prevalence for specific combinations of annotator-target attributes. To ease the exploration of our results, an interactive version of this visualization is available at https://shorturl.at/PC7vK (Abrate, Bacciu, and Cima 2024).
@@
Results. We found statistically significant labeling differences for 17 of 29 (59%) socio-demographic attributes. For each of these cases, we measured bias intensity ($I$), prevalence ($P$), and annotator agreement ($\kappa$). Results are shown in Figure 3. Regarding the 17 significant differences, 11 (65%) were overestimations ($I>0$) and 6 were underestimations. Bias intensity was generally mild or moderate, with $I\leq\pm 0.25$ in most cases. However, we identified strong biases ($I>\pm 0.5$) in some instances. As expected, the strongest biases occur when there is little agreement ($\kappa<0.4$) between the annotator groups. Regarding prevalence $P$, some biases are overall marginal since they occur very infrequently. There are situations, however, for which we found biases that are both intense and with large prevalence ($P>0.4$), such as those related to transgender annotators and targets. Other cases with considerable biases are those annotators and targets in the seniors age group and of Mormon religion, when compared to other annotators.
@@
Table 2: Examples of human biases. For each example we report the number of considered labels (#), and the values of bias intensity ($I$), prevalence ($P$), and agreement ($\kappa$). All reported examples are statistically significant at the 0.01 level, as confirmed by a Mann-Whitney test. Colored symbols map examples into the spaces of Figure 5.
@@
Results. We observed statistically significant labeling differences in 122 out of 1,950 (6%) combinations of annotator and target attributes when using persona-based LLMs. In comparison, human-generated annotations showed significant differences in 1,200 out of 1,950 (62%) combinations. The relatively low number of significant results for persona-based LLMs, especially compared to human annotators, suggests that the personalization had a limited impact on the hate speech labels assigned by the models and that the socio-demographic attributes embedded in the LLMs did not lead to substantial labeling differences. Figure 6 presents each significant combination in the $I$-$\kappa$ space (inset) and $I$-$P$ space (outer figure). Unlike human annotators, persona-based LLMs largely exhibited substantial agreement ($\kappa\geq 0.6$) in their label assignments. In terms of bias intensity, we identified a range from mild to severe, with both notable over- and underestimations. However, the high level of agreement resulted in relatively low bias prevalence, indicating that the majority of detected biases were infrequent. Nevertheless, we identified several biases characterized by both high intensity and substantial prevalence. For instance, LLMs impersonating annotators who identified as Christian consistently underreported hate directed at bisexual individuals ($I=-0.84$, $P=0.27$, $p<0.01$). Similarly, LLMs impersonating annotators with straight sexual orientation underreported hate targeted at men ($I=-0.91$, $P=0.26$, $p<0.01$), while those impersonating bisexual ($I=0.91$, $P=0.27$, $p<0.01$) and gay ($I=0.89$, $P=0.23$, $p<0.01$) annotators overreported it.
@@
Results. We found 86 statistically significant ($p<0.01$) cases in which human annotators and persona-based LLMs exhibited a bias for the same combination of annotator and target attributes. As shown in Figure 7A, the Pearson correlation $r=-0.023$ between the intensity of human and LLM biases suggests the lack of a linear relationship. This near-zero correlation indicates that, overall, the biases exhibited by persona-based LLMs do not correspond to those observed in human annotators. A notable exception is the marked tendency that men annotators—both human and LLM-impersonated—exhibit to overestimate hate towards gay individuals, with bias intensity $I=0.74$ and $I=0.89$ for humans and LLMs respectively. Instead, Figure 7B shows that the correlation between the prevalence of human and LLM biases is moderately positive, with $r=0.388$. This suggests that the biases that are (in)frequent among human annotators tend to be relatively (in)frequent among persona-based LLM annotations as well. Finally, the correlation $r=0.182$ between the agreement measured for human and LLM biases, shown in Figure 7C, indicates a weak positive relationship. This reflects a slight tendency for the level of agreement among human annotators (with and without a certain attribute) to correspond with the level of agreement among LLMs (with and without the same attribute). However, the weak correlation suggests that while there is some degree of similarity in how humans and LLMs agree within and across certain socio-demographic attributes, the factors influencing agreement among humans are not fully captured by LLMs. As already observed in Figure 6, this is mainly due to the large agreement exhibited by the persona-based LLMs, which is indicative of strong labeling consistency even when LLMs are personalized with different attributes.
@@
Bias intensity and prevalence. Our methodology introduces two bias indicators—intensity ($I$) and prevalence ($P$)—that measure the severity and frequency of annotation biases. While $P$ indicates the frequency of disagreement, $I$ measures the extent of imbalance when such disagreement occurs. Clearly, the most concerning biases are those with both high intensity and prevalence. However, even biases with high intensity and small but non-negligible prevalence are attention worthy, as they represent rare but highly skewed annotations that might warrant further investigation.
@@
Human biases. We systematically applied our methodology to identify and characterize human biases in a hate speech annotation task, uncovering a large number of biases with varying intensity and prevalence. Therefore our study favors breadth over depth, and it is beyond our scope to investigate the identified biases in detail. However, our work provides a foundation for domain experts to conduct in-depth studies on specific biases, exploring their origins and potential solutions. To foster this line of work, we publicly release the complete list of statistically significant biases that we found, and the indicators that characterize them. [^7] https://doi.org/10.5281/zenodo.13906773
@@
LLM biases. We also investigated the biases exhibited by persona-based LLMs, and compared them to human ones. Current research in this area is facing a significant tension: on one hand, we strive for unbiased LLMs to avoid propagating human prejudices (Perez et al. 2023). On the other hand, there is a compelling case for developing LLMs that faithfully reproduce human behaviors—including biases—as this would open new frontiers in social simulations (Törnberg et al. 2023). Our results in this regard are concerning. First, we found that persona-based LLMs obtained via prompting do exhibit biases. More troubling, however, is the finding that there is very little alignment between the biases exhibited by LLMs and those exhibited by humans. This suggests that current persona-based LLMs may fail to accurately reflect the complex and nuanced biases present in human behavior. This raises concerns about the reliability of using prompting techniques to create persona-based LLMs for social simulations and underscores the need for further research into the sources of LLM biases.
@@
LLM personalization. Our study is the first to analyze bias in hate speech annotations at such a fine-grained level. Thus, our results about the inability of persona-based LLMs to reproduce human biases might stem from two factors: the LLMs’ inability to replicate fine-grained biases, or the ineffectiveness of the personalization process. About the former, more powerful LLMs than those tested herein could potentially overcome the issue, resulting in more accurate human representations. Instead, regarding personalization, we found a limited capacity of the LLMs to modify their behavior and adapt to specific personas. This result reinforces the current body of work that questions the ability of LLMs to faithfully reproduce human behaviors (Santurkar et al. 2023; Lee, An, and Thorne 2023). Nonetheless, although we used a widely adopted state-of-the-art approach for personalization (Beck et al. 2024), different personalization methods (e.g., model fine-tuning) might yield better results.

### v4

@@ Abstract
[…]
Warning: This paper contains examples that may be perceived as offensive or upsetting. Reader discretion is advised.
@@ Introduction
The proliferation of hateful and toxic speech on online social platforms is among the most severe threats to inclusive online spaces (Mathew et al. 2019). The exponential increase in user-generated content has made manual hate speech detection and mitigation impractical, driving the development of automated systems (Garg et al. 2023; Cima et al. 2024). However, the effectiveness of automated systems is heavily contingent upon the quality of their training data, which is typically human-labeled. Challenges arise as hate speech detection is inherently subjective: what one individual may perceive as offensive, another might consider acceptable or even normal. This subjectivity is shaped by the annotators’ backgrounds, experiences, and cultural contexts. As such, socio-demographic attributes such as age, gender, race, and others, can influence how annotators interpret and label content, possibly leading to biases (Davani et al. 2023). In turn, biases may seep into the automated detection systems trained on such data, causing severe consequences such as the unwarranted removal of non-hateful content or the disproportionate targeting of certain categories of users (Nogara et al. 2025).
@@
Large language models (LLMs) have recently extended the set of AI tools in support of platform administrators. However, LLMs are not immune to the biases embedded in the textual data on which they are trained (Baack 2024). Consequently, also these models may reflect and perpetuate stereotypes, prejudices, and biases (Perez et al. 2023). Despite the development of various techniques aimed at reducing LLM biases, this remains an outstanding challenge (Ouyang et al. 2022). Interestingly, the capacity of LLMs to produce subjective outputs can also serve beneficial purposes. LLMs can be personalized and deployed as actors in agent-based simulations of online human interactions (Törnberg et al. 2023; Park et al. 2024; Rossetti et al. 2024), which represents a promising way to compensate for the limited availability of platform data in the post-API age. By accurately reproducing certain human biases, persona-based LLMs could lay the groundwork for the next generation of AI-driven social simulations. For these reasons, investigating biases in both human hate speech annotations and persona-based LLMs is a task of great theoretical and practical relevance (Tseng et al. 2024).
@@
In RQ1a, we hypothesize that identification with the target group (i.e., the annotator’s in-group) may heighten the annotator’s sensitivity to hate, leading them to flag such content more often (Sachdeva et al. 2022b; Hu et al. 2024). Conversely, when the target belongs to an out-group, the annotator’s perception of the hate may be less acute. RQ1b broadens the analysis by examining whether certain groups systematically label content differently than others, in relation to the targets of the hate. This analysis reveals possible disparities in how hate speech is perceived across diverse socio-demographic backgrounds.
@@
Bias in human hate speech annotations. The impact that biased datasets have on the reliability and fairness of automated hate speech detection systems sparked much research on the issue (Garg et al. 2023). Waseem and Hovy (2016) showed that cultural and socio-economic backgrounds lead to variability in hate speech annotations. Inherent prejudices further skew labels, as human annotators may be more sensitive to hate speech against familiar groups while underestimating it against less familiar ones. Others found that socio-demographic factors such as age, gender, race and ethnicity, education, and English proficiency, influence annotation outcomes, with younger and minority annotators more likely to label content as hate speech (Al Kuwatly, Wich, and Groh 2020; Sap et al. 2022; Davani et al. 2023). Additionally, annotator beliefs and demographics can introduce further inconsistencies in crowdsourced annotations (Hettiachchi et al. 2023). Although much research has explored socio-demographic biases in hate speech annotations (Garg et al. 2023), most studies focused on a limited set of attributes, primarily race and ethnicity, gender, and age. Moreover, none investigated the interplay between annotator and target attributes. Similar results and limitations surface from existing works on demographic annotation bias for tasks different than hate speech detection, such as sentiment analysis (Ding et al. 2022); speech recognition (Meyer et al. 2020); image annotation (Haliburton et al. 2024); as well as Q&A, text rewriting, offensiveness and politeness rating (Pei and Jurgens 2023).
@@
Demographic alignment of LLMs. LLM prompting has been used to experiment with persona-based and role-playing models, to test the alignment of LLM predictions with human opinions (Tseng et al. 2024). Studies involved replicating famous cognitive and social experiments (Aher, Arriaga, and Kalai 2023; Srivastava et al. 2023), as well as assessing agreement and personalization capabilities. Among the existing works, Beck et al. (2024) and Hu and Collier (2024) measured the impact of socio-demographic prompting on model performance, showing that it can enhance zero-shot learning in subjective tasks, while Schäfer et al. (2024) assessed alignment with human annotations. Additional studies explored the personalities (e.g., Big5, MBTI) exhibited by LLMs when prompted to emulate certain human socio-psychological traits (Rao, Leung, and Miao 2023; La Cava, Costa, and Tagarelli 2024). Results showed marked differences between models and a moderate predisposition to personalization and psychological prompting. However, Santurkar et al. (2023) measured the alignment between human and LLM opinions across a wide array of demographic groups and topics, finding substantial misalignment. The misalignment persisted even after explicitly steering the LLMs towards particular groups. Similarly, Lee, An, and Thorne (2023) investigated the alignment of instruction-tuned LLMs with the disagreement among human annotators, finding that models fail to capture human disagreements. As briefly summarized, there are many contrasting results on the alignment of persona-based LLMs with human opinions. This gap in understanding underscores the need for further research to clarify the extent and conditions under which LLMs mirror human perspectives.
@@
The dataset follows the FAIR principles. It is available on a prominent cloud storage service, [^1] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech (accessed: 15/12/2024) making it findable and accessible. It is interoperable in that it is released in a machine-readable format, and reusable due to the information contained in the original paper (Sachdeva et al. 2022a).
@@
Operationalization. Our definition of bias entails the systematic comparison between the labels assigned by human or LLM annotators with certain attributes, with the labels assigned by annotators without such attributes. We operationalize the comparison by extending the method of Wich et al. (2021) as follows. Let $A=\{a_{1},\ldots,a_{n}\}$ be the set of annotators and $G=\{g_{1},\ldots,g_{m}\}$ be the set of targets. We first fix one socio-demographic attribute $\mathbf{t}$ and a specific value of that attribute $\mathbf{t}=v$ (e.g., ideology $=$ liberal). Next, we select all posts labeled by all annotators $A(\mathbf{t}=v)$ with the fixed attribute and value. Depending on the RQ to answer, we further filter the posts by only selecting those whose target $G(\mathbf{t^{\prime}}=v^{\prime})$ feature another, possibly different, attribute and value $\mathbf{t^{\prime}}=v^{\prime}$. Finally, we compare the labels assigned by the annotators $A(\mathbf{t}=v)$ to those of all other annotators $A(\mathbf{t}\neq v)$ on the same set of selected posts. This approach to quantifying annotation bias measures differences between the labels assigned by two groups of annotators, without requiring a fixed baseline or a ground-truth. In other words, it captures relative labeling differences between groups without prioritizing one group’s evaluations over another. The comparison is based on the confusion matrix $D$ in Figure 2. Rows in $D$ correspond to the labels assigned by the annotators $A(\mathbf{t}=v)$, while columns correspond to the labels assigned by $A(\mathbf{t}\neq v)$. The diagonal represents those instances that are labeled equally by the two groups. The lower triangle represents instances where the annotators $A(\mathbf{t}=v)$ overestimate the hate with respect to $A(\mathbf{t}\neq v)$. Conversely, the upper triangle represents instances where $A(\mathbf{t}=v)$ underestimate hate. We repeat the analysis $\forall\;\mathbf{t},v$. Algorithm A1 in the Appendix provides the pseudo-code implementation of the process used to compute the confusion matrix $D$.
Statistical significance. Our comprehensive approach to investigating bias across a large set of socio-demographic attributes results in a substantial number of comparisons. However, as reported in Appendix Table T1, certain minority groups of annotators and targets are underrepresented in the data. Consequently, some comparisons are based on limited data, and the observed labeling differences may lack statistical significance. We address this issue by conducting Mann-Whitney tests to assess the statistical significance of the differences between the distributions of labels assigned by any two groups of annotators being compared. Additionally, given the large number of comparisons, we apply the Holm-Bonferroni correction for multiple hypothesis testing (Abdi 2010). In the remainder, we only report results about those comparisons whose p-value $<0.1$. Due to data imbalance, we were unable to obtain any significant result for a small set of particularly underrepresented groups, highlighted in gray in Table T1.
@@
Model selection. We prioritize open-source models over commercial options (e.g., GPT-4) when selecting the LLMs for our experiments. First, open-source models are freely available, which enhances transparency and reproducibility of our research. Additionally, many open-source models have comparable accuracy to that of leading commercial LLMs. [^2] https://lmarena.ai/?leaderboard (accessed: 15/12/2024) Open-source LLMs also require less computational resources, broadening accessibility and easing implementation. These choices make our results generalizable and broadly applicable. Driven by these considerations, we based our model selection on recent results on the performance of open-source LLMs, favoring high-performance and widely-used models (La Cava, Costa, and Tagarelli 2024). Our selection includes Llama-3-8B-Instruct (8B), [^3] https://shorturl.at/0xT0b (accessed: 15/12/2024) Phi-3-mini-4k-instruct (3.8B), [^4] https://shorturl.at/tOs9i (accessed: 15/12/2024) SOLAR-10.7B-Instruct-v1.0 (10B), [^5] https://shorturl.at/udLQZ (accessed: 15/12/2024) and Starling-LM-7B-alpha (7B). [^6] https://shorturl.at/kqqLq (accessed: 15/12/2024)
@@
Figure 4: Intensity ($I$) and prevalence ($P$) of human biases in hate speech annotations, for all possible combinations of annotator (y axis) and target (x axis) socio-demographic attributes. Cell color corresponds to intensity while cell area is proportional to prevalence. Statistical significance is assessed at the 0.1 level with a Mann-Whitney test with Holm-Bonferroni correction. Missing cells denote non-significant results. Red-colored cells mark statistically significant hate overestimations, while blue-colored ones denote significant underestimations. The magenta-colored top-left square corresponds to the socio-demographic attributes that are available for both the annotators and targets (age, gender, race, religion, sexuality), as thoroughly reported in Appendix Table T1. The dashed diagonal corresponds to in-group annotations (i.e., same annotator and target attribute). Attributes outside of the top-left square are only available either for annotators (education, ideology, income) or targets (disability, origin). Annotator and target attributes for which we did not obtain any significant result (e.g., ideology: extremely liberal) are not shown. When observed row-wise, the heatmap allows identifying groups of annotators with the tendency to systematically over- or underestimate hate. When observed column-wise, it allows identifying hate targets that are overall flagged more or less frequently. Single cells show bias intensity and prevalence for specific combinations of annotator-target attributes. To ease the exploration of our results, an interactive version of this visualization by Abrate, Bacciu, and Cima (2024) is available at https://shorturl.at/PC7vK.
@@
Results. We found statistically significant labeling differences for 16 of 29 (55%) socio-demographic attributes. For each of these cases, we measured bias intensity ($I$), prevalence ($P$), and annotator agreement ($\kappa$). Results are shown in Figure 3. Regarding the 16 significant differences, 10 (62%) were overestimations ($I>0$) and 6 were underestimations. Bias intensity was generally mild or moderate, with $I\leq\pm 0.25$ in most cases. However, we identified strong biases ($I>\pm 0.5$) in some instances. As expected, the strongest biases occur when there is little agreement ($\kappa<0.4$) between the annotator groups. Regarding prevalence $P$, some biases are overall marginal since they occur very infrequently. There are situations, however, for which we found biases that are both intense and with large prevalence ($P>0.4$), such as those related to transgender annotators and targets. Other cases with considerable biases are those annotators and targets in the seniors age group and of Mormon religion, when compared to other annotators.
@@
Table 2: Examples of human biases. For each example we report the number of considered labels (#), and the values of bias intensity ($I$), prevalence ($P$), and agreement ($\kappa$). All reported examples are statistically significant at the 0.01 level, as confirmed by a Mann-Whitney test with Holm-Bonferroni correction. Colored symbols map examples into the spaces of Figure 5.
@@
Results. We observed statistically significant labeling differences in 65 out of 1,950 (3.33%) combinations of annotator and target attributes when using persona-based LLMs. In comparison, human-generated annotations showed significant differences in 900 out of 1,950 (46%) combinations. The relatively low number of significant results for persona-based LLMs, especially compared to human annotators, suggests that the personalization had a limited impact on the hate speech labels assigned by the models and that the socio-demographic attributes embedded in the LLMs did not lead to substantial labeling differences. Figure 6 presents each significant combination in the $I$-$\kappa$ space (inset) and $I$-$P$ space (outer figure). Unlike human annotators, persona-based LLMs largely exhibited substantial agreement ($\kappa\geq 0.6$) in their label assignments. In terms of bias intensity, we identified a range from mild to severe, with both notable over- and underestimations. However, the high level of agreement resulted in relatively low bias prevalence, indicating that the majority of detected biases were infrequent. Nevertheless, we identified several biases characterized by both high intensity and substantial prevalence. For instance, LLMs impersonating annotators who identified as Christian consistently underreported hate directed at bisexual individuals ($I=-0.84$, $P=0.27$, $p<0.01$). Similarly, LLMs impersonating annotators with straight sexual orientation underreported hate targeted at men ($I=-0.91$, $P=0.26$, $p<0.01$), while those impersonating bisexual ($I=0.91$, $P=0.27$, $p<0.01$) and gay ($I=0.89$, $P=0.23$, $p<0.01$) annotators overreported it.
@@
Results. We found 36 statistically significant ($p<0.1$) cases in which human annotators and persona-based LLMs exhibited a bias for the same combination of annotator and target attributes. As shown in Figure 7A, the Pearson correlation $r=-0.105$ between the intensity of human and LLM biases suggests the lack of a linear relationship. This near-zero correlation indicates that, overall, the biases exhibited by persona-based LLMs do not correspond to those observed in human annotators. A notable exception is the marked tendency that men annotators—both human and LLM-impersonated—exhibit to overestimate hate towards gay individuals, with bias intensity $I=0.74$ and $I=0.89$ for humans and LLMs respectively. Instead, Figure 7B shows that the correlation between the prevalence of human and LLM biases is moderately strong, with $r=0.491$. This suggests that the biases that are (in)frequent among human annotators tend to be relatively (in)frequent among persona-based LLM annotations as well. Finally, the correlation $r=0.124$ between the agreement measured for human and LLM biases, shown in Figure 7C, indicates a weak positive relationship. This reflects a slight tendency for the level of agreement among human annotators (with and without a certain attribute) to correspond with the level of agreement among LLMs (with and without the same attribute). However, the weak correlation suggests that while there is some degree of similarity in how humans and LLMs agree within and across certain socio-demographic attributes, the factors influencing agreement among humans are not fully captured by LLMs. As already observed in Figure 6, this is mainly due to the large agreement exhibited by the persona-based LLMs, which is indicative of strong labeling consistency even when LLMs are personalized with different attributes.
@@
Bias intensity and prevalence. Our methodology introduces two bias indicators—intensity ($I$) and prevalence ($P$)—that measure the severity and frequency of annotation biases. While $P$ indicates the frequency of disagreement, $I$ measures the extent of imbalance when such disagreement occurs. Clearly, the most concerning biases are those with both high intensity and prevalence. However, even biases with high intensity and small but non-negligible prevalence are attention worthy. These correspond to situations where disagreement between groups of annotators occurs overall infrequently, as represented by the low prevalence. However, when disagreements do occur, one group systematically over- or underestimates hate, as reflected by the large intensity. These cases may also warrant further investigation.
@@
Human biases. We systematically applied our methodology to identify and characterize human over- and underestimations in hate speech annotations, uncovering a large number of biases with varying intensity and prevalence. We remark however that we lack a reference or ground-truth value for “correct” hate speech annotations. Therefore, in general, our results should be interpreted as systematic and statistically significant differences between the labels assigned by groups of annotators with different socio-demographic attributes. In other words, while some of the identified labeling differences correspond to negative biases to correct, other instances—such as some overestimations by in-group annotators—might be indicative of genuine insights that annotators with a different background might have missed. To this regard, our study favors breadth over depth, and it is beyond our scope to investigate the identified biases in detail. However, our work provides a foundation for domain experts to conduct in-depth studies on specific biases, exploring their origins and potential solutions, if needed. To foster this line of work, we publicly release the complete list of statistically significant biases that we found, and the indicators that characterize them. [^7] https://doi.org/10.5281/zenodo.13906773
@@
LLM biases. We also investigated the biases exhibited by persona-based LLMs, and compared them to human ones. Current research in this area is facing a significant tension: on one hand, we strive for unbiased LLMs to avoid propagating human prejudices (Perez et al. 2023), especially when these LLMs act as arbiters in content moderation decisions. On the other hand, there is a compelling case for developing LLMs that faithfully reproduce human behaviors—including biases—as this would open new frontiers in social simulations (Törnberg et al. 2023; Park et al. 2024; Rossetti et al. 2024). Our results in this regard are concerning. First, we found that persona-based LLMs obtained via prompting do exhibit biases. More troubling, however, is the finding that there is very little alignment between the biases exhibited by LLMs and those exhibited by humans. This suggests that current persona-based LLMs may fail to accurately reflect the complex and nuanced biases present in human behavior. This raises concerns about the reliability of using prompting techniques to create persona-based LLMs for social simulations and underscores the need for further research into the sources of LLM biases.
@@
LLM personalization. Our study is the first to analyze bias in hate speech annotations at such a fine-grained level. Thus, our results about the inability of persona-based LLMs to reproduce human biases might stem from two factors: the LLMs’ inability to replicate fine-grained biases, or the ineffectiveness of the personalization process. About the former, more powerful LLMs than those tested herein could potentially overcome the issue, resulting in more accurate human representations. Instead, regarding personalization, we found a limited capacity of the LLMs to modify their behavior and adapt to specific personas. This result reinforces the current body of work that questions the ability of LLMs to faithfully reproduce human behaviors (Santurkar et al. 2023; Lee, An, and Thorne 2023). Nonetheless, although we used a widely adopted state-of-the-art approach for personalization (Beck et al. 2024), different personalization methods might yield better results. Among them are approaches based on fine-tuning rather than prompting (Agiza, Mostagir, and Reda 2024), or those relying on specific token sampling functions to induce controlled bias in LLM answers.
@@
Our study also serves as a foundational tool to inform broader discussions on normative, ethical, and practical questions regarding LLM behavior. By providing empirical insights into the biases present in human annotators and persona-based LLMs during hate speech detection, we aim to contribute to the larger discourse. These findings can help guide future research and decision-making in addressing critical issues, such as whether LLMs should aim for unbiased outputs or mirror human decision-making in sensitive contexts. Additionally, our findings have practical implications for platform moderation policies and corporate AI ethics. By identifying socio-demographic biases in hate speech annotation, platforms can refine guidelines and training to ensure fair content moderation. Similarly, corporate AI ethics frameworks can leverage these insights to develop and deploy more equitable AI systems, reducing the risk of perpetuating stereotypes and enhancing fairness in automated decision-making.